# Simple homotopy types of even dimensional manifolds

### v2

@@ 9. Tate cohomology and $\mathbb{Z}C_{2}$-modules
In this section, we will recall some basic facts about $\mathbb{Z}C_{2}$-modules which will be used throughout the proofs of Theorems 11.1, 11.2, LABEL:, and 11.3. We will also explain how the methods of Tate cohomology can be applied to $\mathbb{Z}C_{2}$-modules in preparation for the proof of Theorem 11.3.
@@ 9.1. Tate cohomology
[…]
@@ 9.2. Tate cohomology and the structure of $\mathbb{Z}C_{2}$-modules
Let $A$ be a $\mathbb{Z}C_{2}$-module or, equivalently, an abelian group with an involution ¯$\overline{\cdot}\colon A\to A$. In order to prove Theorems 11.1, 11.2, and 11.3, we would like to find techniques to determine the following groups associated to $A$:
@@
We are especially interested in the case $n$ odd (i.e. $n=1$) but we will consider both cases. Note that, for the groups on the left, the notation $A^{-}=\{x\in A\mid x=-\overline{x}\}$ and $A^{+}=\{x\in A\mid x=\overline{x}\}$ is often used since they are the $(-1)$ and $(+1)$-eigenspaces of the involution action. If $2\in A$ is invertible, then $A\cong A^{+}\oplus A^{-}$ but this need not hold in general.
@@
[…]
@@
[…]
@@
[…]
Remark 9.4.
[…]
The third class of groups can be studied using Tate cohomology due to the following. This is standard (see, for example, [CE56, p. 251]) but we will include a proof here for convenience.
@@
Proof.
Since $C_{2}$ has 2-periodic Tate cohomology group (see, for example, [Br94, VI.9.2]), it suffices to compute $\widehat{H}^{0}(C_{2};A)$ and $\widehat{H}^{-1}(C_{2};A)$. We have $A^{C_{2}}=\{x\in A\mid x=\overline{x}\}$ and $A_{C_{2}}=A/\{x-\overline{x}\mid x\in A\}$ and so the norm map is given by
$N\colon\frac{A}{\{x-\overline{x}\mid x\in A\}}\to\{x\in A\mid x=\overline{x}\},\quad x\mapsto x+\overline{x}.$
The result follows since $\widehat{H}^{-1}(C_{2};A)=\ker(N)$, $\widehat{H}^{0}(C_{2};A)=\operatorname{coker}(N)$. ∎
Remark 9.6.
This shows that, for a group $G$ and $w\colon G\to\{\pm 1\}$ an orientation character, we have
$\mathcal{J}_{n}(G,w)/\mathcal{I}_{n}(G,w)\cong\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))$
which was noted already in Section 2.2.
We now recall a series of special facts about the Tate cohomology of $C_{2}$. The first is a consequence of Proposition 9.2 (i) and the fact that finite cyclic groups have $2$-periodic cohomology (see [Se79, p. 133]). Note that this also applies for $C_{2}$ replaced by an arbitrary finite cyclic group.
@@
[…]
We will next consider results which apply only in the case where $A$ is finite $\mathbb{Z}G$-module, i.e. a $\mathbb{Z}G$-module whose underlying abelian group is finite. This is a consequence of the theory of Herbrand quotients (see [Se79, Chapter VIII, Section 4]) though we will include a direct proof here for convenience.
@@
Proof.
First note that, since $A$ is finite, so is $\widehat{H}^{n}(C_{2};A)$ by Proposition 9.5. Next note that there are exact sequences of finite abelian groups (see, for example, [Se79, p. 134]):
$0\to\{x\in A\mid\overline{x}=x\}\to A\xrightarrow[]{}A\to\frac{A}{\{x-\overline{x}\mid x\in A\}}\to 0$
$0\to\widehat{H}^{1}(C_{2};A)\to\frac{A}{\{x-\overline{x}\mid x\in A\}}\xrightarrow[]{}\{x\in A\mid\overline{x}=x\}\to\widehat{H}^{0}(C_{2};A)\to 0$
where the middle maps are $x\mapsto x+\overline{x}$.
The first implies that |$|\{x\in A\mid\overline{x}=x\}|=|A/\{x-\overline{x}\mid x\in A\}|$ and, by combining this with the second, we get that $|\widehat{H}^{1}(C_{2};A)|=|\widehat{H}^{0}(C_{2};A)|$. Hence, since $\widehat{H}^{n}(C_{2};A)$ is 2-periodic (see Proposition 9.5), we get that $|\widehat{H}^{n}(C_{2};A)|$ is independent of $n\in\mathbb{Z}$.
Next recall that, by Proposition 9.2  (iii), we have $02\cdot\widehat{H}^{n}(C_{2};A)=$ 0, i.e. every element in $\widehat{H}^{n}(C_{2};A)$ has order at most two. This implies that $\widehat{H}^{n}(C_{2};A)\cong(\mathbb{Z}/2)^{d_{n}}$ for some $d_{n}\geq 0$. Since $|\widehat{H}^{n}(C_{2};A)|$ is independent of $n$, this implies that $d_{n}$ is also independent of $n$. ∎
For a finite abelian group $A$ and a prime $p$, let $A_{(p)}=\{x\in A\mid p^{n}\cdot x=0\text{ for some $n\geq 1$}\}$ denote the $p$-primary component of $A$. This coincides with the Sylow $p$-subgroup of $A$ and we use this notation since $A_{(p)}\cong\mathbb{Z}_{(p)}\otimes_{\mathbb{Z}}A$ where $\mathbb{Z}_{(p)}$ denotes the localisation of $\mathbb{Z}$ at $S=\mathbb{Z}\setminus(p)$.
@@
Let $A$ be a finite $\mathbb{Z}C_{2}$-module and let $n\in\mathbb{Z}$. Then
$\widehat{H}^{n}(C_{2};A)\cong\widehat{H}^{n}(C_{2};A_{(2)}).$
In order to prove this, we will use the following standard fact about localisations.
Lemma 9.10.
Let $A$ be a finite $\mathbb{Z}C_{2}$-module. For $p$ prime, $A_{(p)}\leq A$ is a $\mathbb{Z}C_{2}$-submodule and $A\cong\bigoplus_{p\mid|A|}A_{(p)}$ as $\mathbb{Z}C_{2}$-modules.
Proof of Proposition 9.9.
By Lemma 9.10 and Proposition 9.2 (ii), we get that
$\widehat{H}^{n}(C_{2};A)\cong\bigoplus_{p\mid|A|}\widehat{H}^{1}(C_{2};A_{(p)}).$
If $p\neq 2$, then $|A_{(p)}|$ is odd and so $\widehat{H}^{1}(C_{2};A_{(p)})=0$ by Proposition 9.2 (iv). This gives that $\widehat{H}^{n}(C_{2};A)\cong\widehat{H}^{n}(C_{2};A_{(2)})$, as required. ∎
@@ 10. Computing the involution on $\widetilde{K}_{0}(\mathbb{Z}C_{m})$
The aim of this section will be to investigate the involution on $\widetilde{K}_{0}(\mathbb{Z}C_{m})$ in preparation for the proofs of Theorems 11.1, 11.2, LABEL:, and 11.3 in Section 11.
We will view $\widetilde{K}_{0}(\mathbb{Z}C_{m})$ as a $\mathbb{Z}C_{2}$-module with the $C_{2}$-action coming from the standard involution on $\widetilde{K}_{0}$ as defined in Section 5.1. We saw in Section 8.1 that $\widetilde{K}_{0}(\mathbb{Z}C_{m})\cong C(\mathbb{Z}C_{m})$ is an isomorphism of $\mathbb{Z}C_{2}$-modules where $C(\mathbb{Z}C_{m})$ denotes the locally free class group, and in particular it is finite. Our basic approach for computing $C(\mathbb{Z}C_{m})$ will be to use the following short exact sequence of $\mathbb{Z}C_{2}$-modules established in Section 8.4
@@ 10.1. The induced involution on the maximal order
where $\zeta_{d}=e^{2\pi i/d}$ denotes a $d$th primitive root of unity. Since $\mathcal{O}_{\mathbb{Q}(\zeta_{d})}=\mathbb{Z}[\zeta_{d}]$, it follows that $\Gamma_{m}=\prod_{d\mid m}\mathbb{Z}[\zeta_{d}]$ is a maximal order in $\mathbb{Q}C_{m}$. The image of $\mathbb{Z}C_{m}$ under the isomorphism above is contained in $\Gamma_{m}$ and so $\Gamma_{m}$ contains $\mathbb{Z}C_{m}$. In fact, $\Gamma_{m}$ is the unique maximal order in $\mathbb{Q}C_{m}$ containing $\mathbb{Z}C_{m}$ [CR87, p. 243]. This implies that there is an isomorphism of abelian groups
@@
where, as noted in Proposition 8.5, the locally free class group $C(\mathbb{Z}[\zeta_{d}])$ coincides with the ideal class group of $\mathbb{Z}[\zeta_{d}]$.
For an integer $d\geq 1$, let ¯$\overline{\,\cdot\,}\colon C(\mathbb{Z}[\zeta_{d}])\to C(\mathbb{Z}[\zeta_{d}])$ denote the map induced by conjugation, i.e. if $\sigma\colon\mathbb{Z}[\zeta_{d}]\to\mathbb{Z}[\zeta_{d}]$ is the ring homomorphism generated by $\zeta_{d}\mapsto\zeta_{d}^{-1}$, then ¯$\overline{\,\cdot\,}=\sigma_{*}$ is the induced map on $C(\mathbb{Z}[\zeta_{d}])$. We will now compute the induced involution on $\bigoplus_{d\mid m}C(\mathbb{Z}[\zeta_{d}])$. This was shown in the case where $m$ is prime in [Re68] (see also [CR87, p. 275]).
@@
This means that, if $x\in C(\mathbb{Z}C_{m})$ and $i_{*}(x)=\prod_{d\mid m}x_{d}$, then $i_{*}(c(x))=\prod_{d\mid m}c_{d}(x_{d})$ where $c\colon C(\mathbb{Z}C_{m})\to C(\mathbb{Z}C_{m})$ is the standard involution and $c_{d}\colon C(\mathbb{Z}[\zeta_{d}])\to C(\mathbb{Z}[\zeta_{d}])$ is induced by conjugation.
@@
For each $d\mid m$, let $i^{(d)}\colon\mathbb{Z}C_{m}\to\mathbb{Z}[\zeta_{d}]$, $x\mapsto\zeta_{d}$. It suffices to prove that, under the map $i^{(d)}_{*}\colon C(\mathbb{Z}C_{m})\to C(\mathbb{Z}[\zeta_{d}])$, the involution on $C(\mathbb{Z}C_{m})$ induces conjugation on $C(\mathbb{Z}[\zeta_{d}])$. By Proposition 8.13, the standard involution on $C(\mathbb{Z}C_{m})$ is induced by the involution on the idèle group $J(\mathbb{Q}C_{m})$. Note that $i^{(d)}_{*}$ is induced by the map $J(i^{(d)})\colon J(\mathbb{Q}C_{m})\to J(\mathbb{Q}(\zeta_{d}))$. Under the map $i^{(d)}$, the involution on $\mathbb{Q}C_{m}$ induces conjugation on $\mathbb{Q}(\zeta_{d})$. In particular, the involution on $C(\mathbb{Z}[\zeta_{d}])$ induced by $i^{(d)}_{*}$ coincides with the involution induced by conjugation on $J(\mathbb{Q}(\zeta_{d}))$. The result now follows since, if a locally free $\mathbb{Z}[\zeta_{d}]$-ideal $M=(x_{1},\dots,x_{n})\subseteq\mathbb{Q}(\zeta_{d})$ is represented by $\alpha\in J(\mathbb{Q}(\zeta_{d}))$, then $\overline{M}=(\overline{x}_{1},\dots,\overline{x}_{n})\subseteq\mathbb{Q}(\zeta_{d})$ is represented by $\overline{\alpha}\in J(\mathbb{Q}(\zeta_{d}))$. ∎
@@ 10.2. Ideal class groups of cyclotomic fields
For every integer $m\geq 2$, let $\lambda_{m}=\zeta_{m}+\zeta_{m}^{-1}$. Let $i\colon\mathbb{Z}[\lambda_{m}]\hookrightarrow\mathbb{Z}[\zeta_{m}]$ denote inclusion and recall that the map $i_{*}\colon C(\mathbb{Z}[\lambda_{m}])\to C(\mathbb{Z}[\zeta_{m}])$ is injective [La78, Theorem 4.2]. Furthermore, the norm map gives a surjection $N\colon C(\mathbb{Z}[\zeta_{m}])\to C(\mathbb{Z}[\lambda_{m}])$ such that the composition
[…]
is the map $x\mapsto x+\overline{x}$ (see, for example, [La78, pp. 83-4]). By viewing $C(\mathbb{Z}[\zeta_{d}])$ and $C(\mathbb{Z}[\lambda_{m}])$ as $\mathbb{Z}C_{2}$-modules under the conjugation action, the maps $i_{*}$ and $N$ are $\mathbb{Z}C_{2}$-module homomorphisms. Note that the conjugation action induces the identify on $C(\mathbb{Z}[\lambda_{m}])$.
@@
[…]
@@
In order to set up later applications, we will now use Lemma 10.2 to obtain information about each of the following groups:
@@
Since $C(\mathbb{Z}[\zeta_{m}])$ is a finite abelian group (see, for example, Section 8.1), we can define the class number of the cyclotomic integers to be $h_{m}=|C(\mathbb{Z}[\zeta_{m}])|$. By Lemma 10.2 (ii), we have that $h_{m}=h_{m}^{-}h_{m}^{+}$ where $h_{m}^{-}=|C(\mathbb{Z}[\zeta_{m}])^{-}|$ and $h_{m}^{+}=|C(\mathbb{Z}[\lambda_{m}])|$. We refer to $h_{m}^{-}$ as the minus part of the class number and $h_{m}^{+}$ as the plus part of the class number respectively.
For an integer $m$, let $\operatorname{odd}(m)$ denote the odd part of $m$, i.e. $\operatorname{odd}(m)$ is the unique odd integer $r$ such that $m=2^{k}r$ for some $k$.
@@
There are subgroups:
  1. (i)
    $C(\mathbb{Z}[\zeta_{m}])^{-}\leq\{x\in C(\mathbb{Z}[\zeta_{m}])\mid x=-\overline{x}\}$;
  2. (ii)
    $2\cdot C(\mathbb{Z}[\zeta_{m}])^{-}\leq\{x-\overline{x}\mid x\in C(\mathbb{Z}[\zeta_{m}])\}$.
In particular, $h_{m}^{-}$ divides |$|\{x\in C(\mathbb{Z}[\zeta_{m}])\mid x=-\overline{x}\}|$ and $\operatorname{odd}(h_{m}^{-})$ divides |$|\{x-\overline{x}\mid x\in C(\mathbb{Z}[\zeta_{m}])\}|$.
@@
Parts (i) and (ii) each follows from Lemma 9.3, Lemma 10.2 (ii) and the fact that $\overline{x}=-x$ for all $x\in C(\mathbb{Z}[\zeta_{m}])^{-}$. For the last part note that, since $C(\mathbb{Z}[\zeta_{m}])^{-}$ is a finite abelian group, we have $C(\mathbb{Z}[\zeta_{m}])^{-}\cong A\oplus B$ where $|A|$ is even and $|B|=\operatorname{odd}(h_{m}^{-})$ is odd. Since $B$ is odd, $2\cdot B=B$ and so

### v3

@@ 9. Tate cohomology and $\mathbb{Z}C_{2}$-modules
In this section, we recall some basic facts about $\mathbb{Z}C_{2}$-modules which are used throughout the proofs of Theorems 11.1 and 11.2. We also explain how the methods of Tate cohomology can be applied to $\mathbb{Z}C_{2}$-modules in preparation for the proof of Theorem 11.2.
@@ 9.1. Tate cohomology
    Let $0\to A\xrightarrow[]{\alpha}B\xrightarrow[]{\beta}C\to 0$ be a short exact sequence of $\mathbb{Z}G$-modules. Then there is a long exact sequence of Tate cohomology groups:
    $\cdots\to\widehat{H}^{n-1}(G;C)\xrightarrow[]{\partial}\widehat{H}^{n}(G;A)\xrightarrow[]{\alpha_{*}}\widehat{H}^{n}(G;B)\xrightarrow[]{\beta_{*}}\widehat{H}^{n}(G;C)\xrightarrow[]{\partial}\widehat{H}^{n+1}(G;A)\to\cdots$
@@ 9.2. Tate cohomology and the structure of $\mathbb{Z}C_{2}$-modules
Let $A$ be a $\mathbb{Z}C_{2}$-module or, equivalently, an abelian group with an involution $\overline{\cdot}\colon A\to A$. In order to prove Theorems 11.1 and 11.2, we would like to find techniques to determine the following groups associated to $A$:
@@
For the groups on the left, the notation $A^{-}=\{x\in A\mid x=-\overline{x}\}$ and $A^{+}=\{x\in A\mid x=\overline{x}\}$ is often used since they are the (1)(-1) and $(+1)$-eigenspaces of the involution action.
@@
Let $0\to A\xrightarrow[]{\alpha}B\xrightarrow[]{\beta}C\to 0$ be an exact sequence of $\mathbb{Z}C_{2}$-modules and let $n\in\mathbb{Z}$.
@@
    $0\to\{x\in A\mid\overline{x}=(-1)^{n}x\}\xrightarrow[]{\alpha}\{x\in B\mid\overline{x}=(-1)^{n}x\}\xrightarrow[]{\beta}\{x\in C\mid\overline{x}=(-1)^{n}x\}.$
@@
    $\{x+(-1)^{n}\overline{x}\mid x\in A\}\xhookrightarrow{\alpha}\{x+(-1)^{n}\overline{x}\mid x\in B\}\xrightarrow[]{\beta}\mathrel{\mkern-14.0mu}\rightarrow\{x+(-1)^{n}\overline{x}\mid x\in C\}.$
The third class of groups can be studied using Tate cohomology via the following [CE56, p. 251].
@@
We now recall a series of special facts about the Tate cohomology of $C_{2}$. The first is a consequence of Proposition 9.2 (i) and the fact that finite cyclic groups have $2$-periodic cohomology (see [Se79, p. 133]). This also applies for $C_{2}$ replaced by an arbitrary finite cyclic group.
@@
Let $0\to A\xrightarrow[]{\alpha}B\xrightarrow[]{\beta}C\to 0$ be a short exact sequence of $\mathbb{Z}C_{2}$-modules. Then there is a $6$-periodic exact sequence of abelian groups:
$\widehat{H}^{1}(C_{2};A)$$\widehat{H}^{1}(C_{2};B)$$\widehat{H}^{1}(C_{2};C)$$\widehat{H}^{0}(C_{2};C)$$\widehat{H}^{0}(C_{2};B)$$\widehat{H}^{0}(C_{2};A).$$\scriptstyle{\alpha_{*}}$$\scriptstyle{\beta_{*}}\scriptstyle{\partial}\scriptstyle{\partial}$$\scriptstyle{\beta_{*}}$$\scriptstyle{\alpha_{*}}$
The following is a consequence of the theory of Herbrand quotients [Se79, Chapter VIII §4].
@@
For a finite abelian group $A$ and a prime $p$, let $A_{p}=\{x\in A\mid p^{n}\cdot x=0\text{ for some $n\geq 1$}\}$ denote the $p$-primary component of $A$. The following is standard.
@@
Let $A$ be a finite $\mathbb{Z}C_{2}$-module and let $n\in\mathbb{Z}$. Then $\widehat{H}^{n}(C_{2};A)\cong\widehat{H}^{n}(C_{2};A_{2})$.
@@ 10. Computing the involution on $\widetilde{K}_{0}(\mathbb{Z}C_{m})$
The aim of this section is to investigate the involution on $\widetilde{K}_{0}(\mathbb{Z}C_{m})$ in preparation for the proofs of Theorems 11.1 and 11.2 in Section 11. We view $\widetilde{K}_{0}(\mathbb{Z}C_{m})$ as a $\mathbb{Z}C_{2}$-module with the $C_{2}$-action coming from the standard involution on $\widetilde{K}_{0}$ as defined in Section 5.1. We saw in Section 8.1 that $\widetilde{K}_{0}(\mathbb{Z}C_{m})\cong C(\mathbb{Z}C_{m})$ is an isomorphism of $\mathbb{Z}C_{2}$-modules where $C(\mathbb{Z}C_{m})$ denotes the locally free class group, and in particular it is finite. Our basic approach for computing $C(\mathbb{Z}C_{m})$ is to use the following short exact sequence of $\mathbb{Z}C_{2}$-modules established in Section 8.4
@@ 10.1. The induced involution on the maximal order
where $\zeta_{d}=e^{2\pi i/d}$, which is a $d$th primitive root of unity. Since $\mathcal{O}_{\mathbb{Q}(\zeta_{d})}=\mathbb{Z}[\zeta_{d}]$, it follows that $\Gamma_{m}=\prod_{d\mid m}\mathbb{Z}[\zeta_{d}]$ is a maximal order in $\mathbb{Q}C_{m}$. The image of $\mathbb{Z}C_{m}$ under the isomorphism above is contained in $\Gamma_{m}$ and so $\Gamma_{m}$ contains $\mathbb{Z}C_{m}$. In fact, $\Gamma_{m}$ is the unique maximal order in $\mathbb{Q}C_{m}$ containing $\mathbb{Z}C_{m}$ [CR87, p. 243]. This implies that there is an isomorphism of abelian groups
@@
where, by Proposition 8.5, $C(\mathbb{Z}[\zeta_{d}])$ coincides with the ideal class group of $\mathbb{Z}[\zeta_{d}]$.
For an integer $d\geq 1$, let $\overline{\,\cdot\,}\colon C(\mathbb{Z}[\zeta_{d}])\to C(\mathbb{Z}[\zeta_{d}])$ denote the map induced by conjugation, i.e. if $\sigma\colon\mathbb{Z}[\zeta_{d}]\to\mathbb{Z}[\zeta_{d}]$ is the ring homomorphism generated by $\zeta_{d}\mapsto\zeta_{d}^{-1}$, then $\overline{\,\cdot\,}=\sigma_{*}$ is the induced map on $C(\mathbb{Z}[\zeta_{d}])$. We now compute the induced involution on $\bigoplus_{d\mid m}C(\mathbb{Z}[\zeta_{d}])$.
@@
This was shown in the case where $m$ is prime in [Re68] (see also [CR87, p. 275]). Since $\Gamma_{m}$ is a maximal order, the map $i_{*}$ is surjective (see Section 8.2).
@@
For each $d\mid m$, let $i^{(d)}\colon\mathbb{Z}C_{m}\to\mathbb{Z}[\zeta_{d}]$, $x\mapsto\zeta_{d}$ where $x$ denotes the generator of $C_{m}$. It suffices to prove that, under $i^{(d)}_{*}\colon C(\mathbb{Z}C_{m})\to C(\mathbb{Z}[\zeta_{d}])$, the involution on $C(\mathbb{Z}C_{m})$ induces conjugation on $C(\mathbb{Z}[\zeta_{d}])$. By Proposition 8.13, the standard involution on $C(\mathbb{Z}C_{m})$ is induced by the involution on the idèle group $J(\mathbb{Q}C_{m})$. Note that $i^{(d)}_{*}$ is induced by the map $J(i^{(d)})\colon J(\mathbb{Q}C_{m})\to J(\mathbb{Q}(\zeta_{d}))$. Under the map $i^{(d)}$, the involution on $\mathbb{Q}C_{m}$ induces conjugation on $\mathbb{Q}(\zeta_{d})$. In particular, the involution on $C(\mathbb{Z}[\zeta_{d}])$ induced by $i^{(d)}_{*}$ coincides with the involution induced by conjugation on $J(\mathbb{Q}(\zeta_{d}))$. The result now follows since, if a locally free $\mathbb{Z}[\zeta_{d}]$-ideal $M=(x_{1},\dots,x_{n})\subseteq\mathbb{Q}(\zeta_{d})$ is represented by $\alpha\in J(\mathbb{Q}(\zeta_{d}))$, then $\overline{M}=(\overline{x}_{1},\dots,\overline{x}_{n})\subseteq\mathbb{Q}(\zeta_{d})$ is represented by $\overline{\alpha}\in J(\mathbb{Q}(\zeta_{d}))$. ∎
@@ 10.2. Ideal class groups of cyclotomic fields
For every integer $m\geq 2$, let $\lambda_{m}=\zeta_{m}+\zeta_{m}^{-1}$. Let $i\colon\mathbb{Z}[\lambda_{m}]\hookrightarrow\mathbb{Z}[\zeta_{m}]$ denote the inclusion map and recall that $i_{*}\colon C(\mathbb{Z}[\lambda_{m}])\to C(\mathbb{Z}[\zeta_{m}])$ is injective [La78, Theorem 4.2]. Furthermore, the norm map gives a surjection $N\colon C(\mathbb{Z}[\zeta_{m}])\to C(\mathbb{Z}[\lambda_{m}])$ such that the composition
$C(\mathbb{Z}[\zeta_{m}])\xrightarrow[]{N}C(\mathbb{Z}[\lambda_{m}])\xrightarrow[]{i_{*}}C(\mathbb{Z}[\zeta_{m}])$
is the map $x\mapsto x+\overline{x}$ (see, for example, [La78, pp. 83-4]). By viewing $C(\mathbb{Z}[\zeta_{d}])$ and $C(\mathbb{Z}[\lambda_{m}])$ as $\mathbb{Z}C_{2}$-modules under the conjugation action, the maps $i_{*}$ and $N$ are $\mathbb{Z}C_{2}$-module homomorphisms. The conjugation action induces the identify on $C(\mathbb{Z}[\lambda_{m}])$.
@@
    $0\to C(\mathbb{Z}[\zeta_{m}])^{-}\to C(\mathbb{Z}[\zeta_{m}])\xrightarrow[]{N}C(\mathbb{Z}[\lambda_{m}])\to 0.$
@@
In order to set up later applications, we now use Lemma 10.2 to obtain information about each of the following groups:
@@
Since $C(\mathbb{Z}[\zeta_{m}])$ is a finite abelian group (see Section 8.1), we can make the following definition.
Definition 10.4.
Define the class number of the cyclotomic integers to be $h_{m}:=|C(\mathbb{Z}[\zeta_{m}])|$. Define the minus part of the class number to be $h_{m}^{-}:=|C(\mathbb{Z}[\zeta_{m}])^{-}|$ and the plus part of the class number to be $h_{m}^{+}:=|C(\mathbb{Z}[\lambda_{m}])|$. By Lemma 10.2 (ii), we have $h_{m}=h_{m}^{-}h_{m}^{+}$.
For an integer $m$, let $\operatorname{odd}(m)$ denote the unique odd integer $r$ such that $m=2^{k}r$ for some $k$.
@@
$\operatorname{odd}(h_{m}^{-})$ divides $|\{x-\overline{x}\mid x\in C(\mathbb{Z}[\zeta_{m}])\}|$.
@@
Since $\overline{x}=-x$ for all $x\in C(\mathbb{Z}[\zeta_{m}])^{-}$, we have $2\cdot C(\mathbb{Z}[\zeta_{m}])^{-}\leq\{x-\overline{x}\mid x\in C(\mathbb{Z}[\zeta_{m}])\}$. Since $C(\mathbb{Z}[\zeta_{m}])^{-}$ is a finite abelian group, we have $C(\mathbb{Z}[\zeta_{m}])^{-}\cong A\oplus B$ where $|A|$ is even and $|B|=\operatorname{odd}(h_{m}^{-})$ is odd. Since $B$ is odd, $2\cdot B=B$ and so